Determine the polynomial function whose graph passes through the points (0, 1), (2,1/3) and (4,1/5). Also sketch the graph of the polynomial function. 


1
Expert's answer
2021-03-26T08:39:07-0400

Solution.

Each of the points "(0,1), (2,\frac{1}{3}),(4,\frac{1}{5})" satisfies the equation "p=ax^2+bx+c" for some unknown "a, b, c."  Substitute each point in the equation and make a matrix equation:


"AX=B, \text{or}""\begin{pmatrix}
 0 & 0 &1 \\
 4 & 2 & 1\\
16 & 4 & 1
\end{pmatrix}\cdot
\begin{pmatrix}
 a \\
b\\
 c 
\end{pmatrix}
=\begin{pmatrix}
 1 \\
\frac{1}{3}\\
 \frac{1}{5}
\end{pmatrix}"

Solve matrix equation: "X = A^{-1} \u00b7 B."

"X=\begin{pmatrix}
 \frac{1}{15} \\
-\frac{7}{15}\\
 1
\end{pmatrix}," or "a=\frac{1}{15}, b=-\frac{7}{15}, c=1."

So, "p=\frac{1}{15}x^2-\frac{7}{15}x+1." It is the polynomial function.